OpenStudy (anonymous):

On the graph below are three points X, Y, ans Z. Write the coordinates for the given dilation. DO, 3 of X = (7, 3) (12, 0) (1, -3)

OpenStudy (anonymous):

@AlexandervonHumboldt2

OpenStudy (anonymous):

do you have choices for it?

OpenStudy (anonymous):

(7, 3) (12, 0) (1, -3

OpenStudy (perl):

multiply each coordinate of X by 3

OpenStudy (anonymous):

thnx:)

OpenStudy (perl):

X = ( 4,0) 3*X = ( 3*4 , 3*0) =

OpenStudy (anonymous):

hope it helped also i agree with @perl too

OpenStudy (perl):

assuming it is dilating with respect to the center

OpenStudy (anonymous):

12,0